Obtain the geometric representation of
((5+i)(7+6i))/3+i
1
Expert's answer
2020-05-21T17:32:55-0400

((5+i)(7+6i))/(3+i)=(35+30i+7i+6i2)/(3+i) =(35+37i+6(-1))/(3+i) =(29+37i)/(3+i)


(29+37i)/(3+i)*(3-i)/(3-i) =(87-29i+111i-37i2)/(9-i2) =(124+82i)/10 =12.4+8.2i

r=(12.42++8.22)0.5=14.9


"\alpha" =arctan(8.2/12.4)=0.58

0.58 radian=33.2°

Answer: z=14.9(cos(33.2°)+isin(33.2°))